# HueManity: Probing Fine-Grained Visual Perception in MLLMs

### v2

@@ Appendix A A Brief Discussion on CIEDE2000 Color Difference
$\text{where}\quad&\Delta L^{\prime}\text{ is the corrected lightness difference,}\\ &\Delta C^{\prime}\text{ is the corrected chroma difference,}\\ &\Delta H^{\prime}\text{ is the corrected hue difference,}\\ &K_{L},K_{C},K_{H}\text{ are parametric factors (typically 1),}\\ &S_{L},S_{C},S_{H}\text{ are weighting functions for lightness, chroma, and hue,}\\ &R_{T}\text{ is a rotation term accounting for hue-chroma interaction.}\end{aligned}$ (1)
The CIEDE2000 score ($\Delta E_{2000}$, Equation A) Luo et al. (2001) quantifies the perceived difference between two colors more accurately than prior formulae, especially for subtle variations. It calculates a single value representing the “distance” between colors in the perceptually uniform CIE $L^{*}a^{*}b^{*}$ space, considering lightness, chroma, and hue. In the HueManity benchmark, $\Delta E_{2000}$ was pivotal for systematically designing stimuli. The ability to discern characters in the Ishihara-style plates directly depends on the perceived color contrast between foreground (character) and background dots. This score provided a perceptually relevant, objective method to quantify this contrast, enabling the selection of color pairs across a controlled spectrum of difficulty, refer to Figure 3. This ensured the benchmark could rigorously test visual perception for varying degrees of color discriminability while maintaining stimuli legibility for human comparison, forming a foundational aspect of our dataset’s controlled experimental design.
@@
Appendix B Color Pairs Selection
@@
    Initial Candidate Generation: We bootstrapped the process with  15 medium-contrast color pairs generated by LLMs (Gemini, ChatGPT). This initial pool was iteratively refined by evaluating pairs against CIEDE2000 ($\Delta E_{2000}$, Eq. A) scores and visual checks. We retained promising candidates, modified some, and discarded others, while simultaneously manually crafting and vetting new pairs to meet the benchmark’s final requirements (detailed below). This refinement cycle culminated in the selection of 25 distinct pairs for the subsequent validation stages.
@@
    Quantitative Contrast Filtering (CIEDE2000): Each of these candidate pairs then underwent rigorous quantitative analysis using the CIEDE2000 ($\Delta E_{2000}$) color difference formula (Equation A). This formula is a standard measure in color science, designed to reflect perceptually meaningful differences as perceived by humans. We established a specific target range for the $\Delta E_{2000}$ score, retaining only pairs with contrast values between 25 and 75. The lower bound of 25 was set to ensure sufficient theoretical distinguishability for individuals with normal color vision, preventing pairs that would be inherently ambiguous. The upper bound of 75 aimed to exclude pairs with excessively high contrast, which might render the perceptual task trivial and deviate from the subtle challenge intended.
@@
    Balanced Contrast Distribution: A key objective during selection was to ensure the benchmark included stimuli across a spectrum of difficulty levels related to color similarity. Therefore, we deliberately curated the final set of 25 color pairs to achieve an approximately equal distribution around a $\Delta E_{2000}$ score of 50. This threshold is grounded in color science literature, often considered a point distinguishing more subtle (scores <50) from more clearly distinct (scores >50) color differences. We aimed for roughly half the selected pairs to fall below this threshold and half above, ensuring HueManity evaluates performance across varying, literature-informed degrees of color contrast difficulty.
@@
    Manual Verification and Legibility Check: Recognizing that a single numerical contrast score like $\Delta E_{2000}$ captures overall perceived difference but may not fully account for the complex interplay of hue, saturation, and luminance components, especially when rendered as dots and subjected to further transformations (gradient, color, and light shifts as described in Section 3.2), a crucial final step of manual verification was performed. As you noted, it is hard to quantify the nuanced visual impact of these combined factors with a single metric. Therefore, for every color pair that passed the quantitative filtering, sample HueManity images were generated. These renderings were meticulously inspected by the authors. The primary goal was to reject pairs where the characters, despite an acceptable overall contrast score, appeared visually too similar to the background due to the specific combination of hue, saturation, luminance, or the effect of the applied shifts. This ensured that the embedded alphanumeric characters were clearly legible and that the pattern recognition was unambiguous for human observers with normal color vision. Any pairs that resulted in ambiguous characters or were otherwise problematic during this visual check were discarded.
@@
Appendix C Qualitative Analysis of MLLM Failure Patterns
@@
C.1 Prevalent Hallucination of Unrelated or Overly Complex Characters
@@
C.2 Frequent Resort to Descriptive Evasion or Explicit Admission of Inability
@@
C.3 Erratic, Unpredictable, and Systematically Flawed Output Patterns
MLLM outputs were frequently characterized by their erratic and unpredictable nature. This included the generation of seemingly random strings of characters, peculiar systematic but incorrect patterns, or extreme numerical inventions far removed from the two-character target. This high variance in error types was observed both across different models for the same input and within the outputs of a single model across different images. For instance, when presented with the same alphanumeric target (e.g., “Wh”), while one model (GPT-4.1) might respond almost correctly, others exhibited diverse failures: Claude 3.7 Sonnet produced an unrelated number (“4726”); LLaVA-13b generated an exceptionally long string of sequential numbers; and Qwen VL Max incorrectly reasoned the presence of a different number (“12”). Some incorrect outputs also suggested flawed systematic processing, such as LLaVA-13b responding with a patterned string like “[L1L1L1]” for one target or generating extremely long, patterned numeric strings for others. Lengthy, seemingly gibberish character strings were also common from models like LLaVA-7b. This unpredictability underscores a lack of robust and stable visual feature extraction and interpretation, contrasting with human visual processing, which tends towards predictable errors based on similarity.

### v5

@@ Appendix 0.C A Brief Discussion on CIEDE2000 Color Difference
[…]
The CIEDE2000 score ($\Delta E_{2000}$, Equation 0.C) [31] quantifies the perceived difference between two colors more accurately than prior formulae, especially for subtle variations. It calculates a single value representing the “distance” between colors in the perceptually uniform CIE $L^{*}a^{*}b^{*}$ space, considering lightness, chroma, and hue. In the HueManity benchmark, $\Delta E_{2000}$ was pivotal for systematically designing stimuli. The ability to discern characters in the Ishihara-style plates directly depends on the perceived color contrast between foreground (character) and background dots. This score provided a perceptually relevant, objective method to quantify this contrast, enabling the selection of color pairs across a controlled spectrum of difficulty, refer to Figure 3. This ensured the benchmark could rigorously test visual perception for varying degrees of color discriminability while maintaining stimuli legibility for human comparison, forming a foundational aspect of our dataset’s controlled experimental design.
@@
Appendix 0.D Color Pairs Selection
@@
    Initial Candidate Generation: We bootstrapped the process with  15 medium-contrast color pairs generated by LLMs (Gemini, ChatGPT). This initial pool was iteratively refined by evaluating pairs against CIEDE2000 ($\Delta E_{2000}$, Eq. 0.C) scores and visual checks. We retained promising candidates, modified some, and discarded others, while simultaneously manually crafting and vetting new pairs to meet the benchmark’s final requirements (detailed below). This refinement cycle culminated in the selection of 25 distinct pairs for the subsequent validation stages.
@@
    Quantitative Contrast Filtering (CIEDE 2000): Each of these candidate pairs then underwent rigorous quantitative analysis using the CIEDE2000 ($\Delta E_{2000}$) color difference formula (Equation 0.C). This formula is a standard measure in color science, designed to reflect perceptually meaningful differences as perceived by humans. We established a specific target range for the $\Delta E_{2000}$ score, retaining only pairs with contrast values between 25 and 75. The lower bound of 25 was set to ensure sufficient theoretical distinguishability for individuals with normal color vision, preventing pairs that would be inherently ambiguous. The upper bound of 75 aimed to exclude pairs with excessively high contrast, which might render the perceptual task trivial and deviate from the subtle challenge intended.
@@
    Balanced Contrast Distribution: A key objective during selection was to ensure the benchmark included stimuli across a spectrum of difficulty levels related to color similarity. Therefore, we deliberately curated the final set of 25 color pairs to achieve an approximately equal distribution around a $\Delta E_{2000}$ score of 50. This threshold is grounded in color science literature, often considered a point distinguishing more subtle (scores << 50) from more clearly distinct (scores >> 50) color differences. We aimed for roughly half the selected pairs to fall below this threshold and half above, ensuring HueManity evaluates performance across varying, literature-informed degrees of color contrast difficulty.
@@
    Manual Verification and Legibility Check: Recognizing that a single numerical contrast score like $\Delta E_{2000}$ captures overall perceived difference but may not fully account for the complex interplay of hue, saturation, and luminance components, especially when rendered as dots and subjected to further transformations (gradient, color, and light shifts), a crucial final step of manual verification was performed. It is hard to quantify the nuanced visual impact of these combined factors with a single metric, therefore for every color pair that passed the quantitative filtering, sample HueManity images were generated. These renderings were meticulously inspected by the authors. The primary goal was to reject pairs where the characters, despite an acceptable overall contrast score, appeared visually too similar to the background due to the specific combination of hue, saturation, luminance, or the effect of the applied shifts. This ensured that the embedded alphanumeric characters were clearly legible and that the pattern recognition was unambiguous for human observers with normal color vision. Any pairs that resulted in ambiguous characters or were otherwise problematic during this visual check were discarded.
@@
Table 7: The 25 Curated Color Pairs Used in the HueManity Benchmark.
Foreground Color Background Color (Foreground | Background) RGB Values
[…]
Appendix 0.E Qualitative Analysis of MLLM Failure Patterns
@@
0.E.1 Prevalent Hallucination of Unrelated or Overly Complex Characters
@@
0.E.2 Frequent Resort to Descriptive Evasion or Explicit Admission of Inability
@@
0.E.3 Erratic, Unpredictable, and Systematically Flawed Output Patterns
MLLM outputs were frequently characterized by their erratic and unpredictable nature. This included the generation of seemingly random strings of characters, peculiar systematic but incorrect patterns, or extreme numerical inventions far removed from the two-character target. This high variance in error types was observed both across different models for the same input and within the outputs of a single model across different images. For instance, when presented with the same alphanumeric target (e.g., “Wh”), while one model (GPT-4.1) might respond almost correctly, others exhibited diverse failures: Claude 3.7 Sonnet produced an unrelated number (“4726”), LLaVA-13b generated an exceptionally long string of sequential numbers, and Qwen VL Max incorrectly reasoned the presence of a different number (“12”). Some incorrect outputs also suggested flawed systematic processing, such as LLaVA-13b responding with a patterned string like “[L1L1L1]” for one target or generating extremely long, patterned numeric strings for others. Lengthy, seemingly gibberish character strings were also common from models like LLaVA-7b. This unpredictability underscores a lack of robust and stable visual feature extraction and interpretation, contrasting with human visual processing, which tends towards predictable errors based on similarity.
Appendix 0.F Usage of Generative AI tools